find the domain of the composite function f of g. f(x)=x^2+9;g(x)=《qrtx-7
please show work

Answer by stanbon(75887)

find the domain of the composite function f of g. f(x)=x^2+9;g(x)=《qrtx-7
----
f[g(x)]
= f[sqrt(x)-7]
=(sqrt(x)-7)^2+9
= x-14sqrt(x)+49 + 9
= x -sqrt(x)+58
------
Domain: x >= 0
====================
cheers,
Stan H.